OpenStudy (anonymous):

QUICK GEOMETRY QUESTION!! easy! It is asking me to simplify the number in the comment and here are my answer choices. 2 2i -2 -2i

OpenStudy (anonymous):

\[\sqrt{-4}\]

OpenStudy (anonymous):

It needs to be i, correct? since it is a negative number under a square root?

OpenStudy (goformit100):

sqrt{-4} = sqrt{-2*2} = 2i

OpenStudy (anonymous):

Okay! thank you! may i ask, what would make the equation so that it turned out to be a negative number? like option -2i

OpenStudy (pixiedust1):

Yeah, goformit100 is correct. Always remember that two negative numbers multiplied equals a positive number.

OpenStudy (goformit100):

x^2 + 4 = 0